OpenStudy (anonymous):

simplify the rational expression state any excluded values x^2-x-6/x+2

OpenStudy (dtan5457):

Factor the trinomial and it becomes another easy one :)

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

I really dont get it. Its my nephews algebra. I have no idea what to do.

OpenStudy (dtan5457):

x^2-x-6 To factor trinomials (this is worth remembering), find 2 numbers that equal -1 when added and -6 when being multiplied

OpenStudy (anonymous):

Still confused

OpenStudy (dtan5457):

Quadratics are set up as ax^2+bx+c as long as the value of a is one, you can say that the answer or roots for a quadratic. Is 2 numbers that add up to the value of b, and multiply to the value of c. x^2-x-6 Think. What two numbers add up to -1 but multiply to -6 (together)

OpenStudy (anonymous):

Still cant get it. Im sorry im just dumb.

OpenStudy (dtan5457):

Don't say that. Example: x^2+4x+4 I have to find two numbers that add up to 4, but multiply to 4 as well. That is simply 2,2 as the numbers (2,2) add up to 4 and multiply to 4.

OpenStudy (dtan5457):

Try one more time

OpenStudy (dtan5457):

You can do it

OpenStudy (anonymous):

Im not understanding what your saying.

OpenStudy (anonymous):

Wait... -2?

OpenStudy (dtan5457):

If I asked you for two numbers that add up to 10, but multiply to 25, could you do it?

OpenStudy (anonymous):

5?

OpenStudy (dtan5457):

Exactly, 5 and 5.

OpenStudy (dtan5457):

x^2-x-6 What two numbers add up to -1 but multiply to -6 (together) Same logic, don't make it hard

OpenStudy (anonymous):

-3

OpenStudy (dtan5457):

That's one of the numbers, what's the other one?

OpenStudy (anonymous):

3

OpenStudy (dtan5457):

two numbers add up to -1 but multiply to -6 3+(-3)=0 3 x -3=-9 that wouldn't work..

OpenStudy (anonymous):

Wait.... no -2

OpenStudy (dtan5457):

Let's see -3,-2 -3+-2=-5 -3 x -2=6 There is a small error with -2

OpenStudy (anonymous):

So just 2

OpenStudy (dtan5457):

Yes!! -3,2 are the factors of x^-x-6 Which can be expressed as (x-3)(x+2)

OpenStudy (dtan5457):

Now onto your question..\[\frac{ (x-3)(x+2) }{ (x+2) }\]

OpenStudy (dtan5457):

What terms can you cross out?

OpenStudy (anonymous):

X+2

OpenStudy (dtan5457):

Exactly so your just left with x-3, the simplified rational expression of your question

OpenStudy (dtan5457):

That took a while but I hope you have a better hint on quadratic factoring now. It's vital in algebra :O

OpenStudy (anonymous):

So the answer is just x-3

OpenStudy (dtan5457):

yep